OpenStudy (anonymous):

x^2-4x-21 need i factored completely

OpenStudy (anonymous):

\[f(-3,7)\] x^2+3x)=7x-21) x(x+3)+7(x-3) (x+7)(x-3)

OpenStudy (precal):

@pastory your factoring is incorrect because it produces x^2+4x-21

OpenStudy (precal):

(x-7)(x+3) check it by using foil x*x=x^2 x*3=3x -7*x=-7x 3x-7x=-4x -7*3=-21